Problem 3 The USDA collects data on commodity prices and consumption. Use their data to explain per-capita chicken consumption as a function of the price of chicken. (Note that economic theory suggests that an increase in a commodity price is generally associated with a decrease in demand.) Q capita chicken consumption PCHICKconsumer price index for chicken We estimated a linear regression model usng annual data for 1950-2001 ( N = 52 ). Dependent variable: coefficient d. error const 0.05 R-squared: 0.854880 a. Write the equation for the least-squares regression line: b.The nu hypothesis is: β Calculate the I-ratio d. The critical value, evaluated at α. 0.05,forthe r distribution (with N-2 df)s . Can we reject the null hypothess? ck ne
Problem 3 The USDA collects data on commodity prices and consumption, Use their data to explain per-capita chicken consumption as a function of the price of chicken. (Note that economic theory suggests that an increase in a commodity price is generally associated with a decrease in demand.) Q per capita chicken consumption PCHICK consumer price index for chicken We estimated a linear regression model using annual data for 1950-2001 ( N = 52 ). Dependent variable: coefficient std.error const 0.05 R-squared: 0.854880 a. Wite the equation for the least-squares regression line: b. The nun hypothesis is: β -Dooren 0 c. Calculate the t-ratio d The critical value, evaluated at α-005, for the t distribution (with N-2 d)s e. Can we reject the null hypothesis?lpick one) yes
Answer #1

a) yhat=5+(-0.1)*x

b) null hypothesis is \beta >=0

c) t -ratio =-0.1/0.05 =-2

d) critical value =-1.676

e) Yes